OpenStudy (anonymous):

Simplify. Express the product as a rational exponent (x^3/5)^2/3

OpenStudy (anonymous):

you multiply the exponents

OpenStudy (anonymous):

3/5 * 2/3 = 6/15

OpenStudy (anonymous):

I got that answer but the choices are x 2/5 , x1/15 , x 19/15 , x 5/8 @hellokitty1997

OpenStudy (anonymous):

simplify 6/15 both the numerator and denominator are divisible by 3... 6/15 = 2/5